# DGInStyle: Domain-Generalizable Semantic Segmentation with Image Diffusion Models and Stylized Semantic Control

### 05 Dec 2023 v1

@@ 4.4 Ablation Studies
MRLF Style Swap Style Prompts RCG Avg3 Avg5
@@ 5 Conclusion
This paper delves into the potential of generative data augmentation using pretrained LDMs in the challenging context of domain generalization for semantic segmentation. We propose DGInStyle, an efficient data generation pipeline that crafts diverse task-specific images by sampling the rich prior of a pretrained latent diffusion model, while ensuring precise adherence of the generation to semantic layout condition. DGInStyle has demonstrated its capability to enhance the generalizability of semantic segmentation models through extensive experiments across various domains. It consistently improves the performance of several domain generalization methods, notably enhancing the state of the art by +2.5 mIoU.

### 08 Apr 2024 v2

@@ 4.4 Ablation Studies
MRLF Swap Prompts RCG Avg3 Avg5
@@ 5 Conclusion
We have explored the potential of generative data augmentation using pretrained LDMs in the challenging context of domain generalization for semantic segmentation. We propose DGInStyle, a novel and efficient data generation pipeline that crafts diverse task-specific images by sampling the rich prior of a pretrained latent diffusion model, while ensuring precise adherence of the generation to semantic layout condition. DGInStyle has demonstrated its capability to enhance the generalizability of semantic segmentation models through extensive experiments across various domains. It consistently improves the performance of several domain generalization methods for both CNN and Transformer architectures, notably enhancing the state of the art. Newly demonstrating the power of LDMs as data generators for domain-robust segmentation, DGInStyle is one more step towards domain-independent semantic segmentation. We hope that it can lay the foundation for future work on how to best utilize generative models for improving domain generalization of dense scene understanding.